# A Study On The Graph Formulation Of Union Closed Conjecture

### v2

@@ 3 Some New Graph Results
Also note that $D$ is a stable set, and since $f^{-1}(D)$ is a subset of $D$, it is also a stable set. Suppose, for contradiction, that $f^{-1}(D)$ is not maximal stable in $C\setminus\Gamma$. This implies there exists a vertex $v\in V(C\setminus\Gamma)$, with $v\notin f^{-1}(D)$, such that the set $\{v\}\cup f^{-1}(D)$ is stable in $C\setminus\Gamma$. Since all vertices in $\Gamma$ are 2-layered, each vertex in $N_{\scriptscriptstyle C}(\Gamma)$ has an adjacent pendant vertex. Due to the maximality of the stable sets, these pendant vertices must be included in all member sets of $\mathcal{B}_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$. Consequently, they are present in every member set of $f^{-1}(\mathcal{B}_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma}))$. Thus, $v\notin N_{\scriptscriptstyle C}(\Gamma)$. Therefore, $\{v\}\cup D=\{v\}\cup f^{-1}(D)\cup\Gamma$ would be stable in $C$, which contradicts the assumption that $D$ is a maximal stable set in $C$. Hence, $f^{-1}(D)\in\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})$ and the mapping is well-defined.
I It is trivial that $f$ is a bijective map. The one-to-one correspondence between the families $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})$ and $\mathcal{B}_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ establishes Equation 1. Similarly, Equation 2 follows from analogous arguments, as it involves collecting all maximal stable sets containing the specified vertex $b\in V(C\setminus\Gamma)$ from both families. ∎
@@
For any $\Theta\subseteq[n]$, $\Gamma_{\scriptscriptstyle 1},\Gamma_{\scriptscriptstyle 2}\subseteq\Theta^{\mathsf{c}}$ and $b\in V(C)$, if $\Gamma_{\scriptscriptstyle 1}\neq\Gamma_{\scriptscriptstyle 2}$ then,
$\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 1}}))\cap\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 2}}))=\phi$ (3)
$\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 1}}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 1}}))\cap\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 2}}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 2}}))=\phi$ (4)
@@
Consider two distinct subsets $\Gamma_{\scriptscriptstyle 1}$ and $\Gamma_{\scriptscriptstyle 2}$ of $\Theta^{\mathsf{c}}$. Without loss of generality, assume that there exists an element $m\in\Gamma_{\scriptscriptstyle 1}$ such that $m\notin\Gamma_{\scriptscriptstyle 2}$. Suppose $B$ is an element of both $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 1})})$ and $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 2})})$. Since $m\in\Gamma_{\scriptscriptstyle 1}$, the maximal stable sets in $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 1})})$ avoid all vertices in $N_{\scriptscriptstyle C}(m)$. Therefore, $B\cap N_{\scriptscriptstyle C}(m)=\emptyset$. On the other hand, since $B$ is also an element of $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 2})})$, it must be a maximal stable set in $C\setminus\Gamma_{\scriptscriptstyle 2}$ that excludes vertices from $\Theta^{\mathsf{c}}$. Given that $m\in\Theta^{\mathsf{c}}$, the maximality of $B$ implies that $B\cap N_{\scriptscriptstyle C}(m)\neq\emptyset$.
This contradiction shows that $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 1})})\cap\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 2})})$ must be empty. Furthermore, it follows that $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})\subseteq\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})$. Consequently, we deduce that $\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 1}}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 1})})\cap\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma_{\scriptscriptstyle 2}}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma_{\scriptscriptstyle 2})})=\emptyset$. ∎
@@
Let $b\in V(C\setminus[n])$ and $a\in[n]$, then :
$w_{\scriptscriptstyle G}=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (5)
$w_{\scriptscriptstyle G}(b)=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle C\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (6)
$w_{\scriptscriptstyle G}(a)=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle C\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (7)
@@
[…]
@@
For simplicity, let us denote the family on the right-hand side by $\mathcal{D}$. Consider any member set $B_{1}\cup B_{2}$ in $\mathcal{D}$. It is straightforward to observe that $B_{1}\cup B_{2}$ forms a stable set in $G$. If this set is not maximal in $G$, then there must exist a vertex $v\in V(G)$ such that $\{v\}\cup B_{1}\cup B_{2}$ is also a stable set. However, due to the maximality of $B_{1}$ and $B_{2}$ in the subgraphs $C\setminus\Gamma$ and $H\setminus\Psi$, respectively, the vertex $v$ cannot belong to either of these subgraphs. Since $C\setminus\Gamma\cup H\setminus\Psi=G$, there is no such vertex $v$ in the graph $G$. Therefore, $B_{1}\cup B_{2}$ must be a maximal stable set in $G$, implying that $B_{1}\cup B_{2}\in\mathcal{B}_{\scriptscriptstyle G}$ and hence $\mathcal{D}\subseteq\mathcal{B}_{\scriptscriptstyle G}$.
@@
Now, consider any $B\in\mathcal{B}_{\scriptscriptstyle G}$. Define the subsets $\Theta\subseteq[n]$ and $\Theta^{\mathsf{c}}=[n]\setminus\Theta$ by setting $\Theta=B\cap[n]$. Next, define $\Gamma,\Psi\subseteq\Theta^{\mathsf{c}}$ as the sets of vertices in $[n]$ such that $N_{\scriptscriptstyle C}(\Gamma)\cap B=\emptyset$ and $N_{\scriptscriptstyle H}(\Psi)\cap B=\emptyset$, respectively. It is evident that $\Gamma\cap\Psi=\emptyset$, because if there were a vertex $v\in\Theta^{\mathsf{c}}$ with $N_{\scriptscriptstyle G}(v)=N_{\scriptscriptstyle C}(v)\cup N_{\scriptscriptstyle H}(v)=\emptyset$, this would contradict the maximality of $B$ as a stable set [^4] Since $v$ and none of its neighbors would be included in $B$. Furthermore, $B\cap(C\setminus\Gamma)$ contains all vertices in $\Theta$ but excludes vertices from both $\Theta^{\mathsf{c}}$ and $N_{\scriptscriptstyle C}(\Gamma)$. Notably, this set is stable since it is a subset of $B$.
@@
Assume that $B$ is not maximal in $C\setminus\Gamma$, and there exists a vertex $v\in V(C\setminus\Gamma)$ such that $v\notin B$ and ($(\{v\}\cup B)\cap(C\setminus\Gamma)$ is a stable set. For each $u\in\Theta^{\mathsf{c}}\setminus\Gamma$, by definition, $N_{\scriptscriptstyle C}(u)\cap B\neq\emptyset$, which implies that $v$ cannot belong to $\Theta^{\mathsf{c}}\setminus\Gamma$. Consequently, $v\notin[n]$. Note that $v$ also does not belong to $N_{\scriptscriptstyle C}(\Theta)$, as ($(\{v\}\cup B)\cap(C\setminus\Gamma)$ remains stable. From this, one can observe that ($(\{v\}\cup B)\cap(H\setminus\Psi)$ is also stable, and consequently, $B\cup\{v\}=((\{v\}\cup B)\cap(C\setminus\Gamma))\cup((\{v\}\cup B)\cap(H\setminus\Psi))$ forms a stable set in $G$, which contradicts the maximality of $B$. Therefore, $B\cap(C\setminus\Gamma)\in\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})$. Similarly, one can show that $B\cap(H\setminus\Psi)\in\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Psi)})$, and hence $B\in\mathcal{D}$. This implies that $\mathcal{B}_{\scriptscriptstyle G}\subseteq\mathcal{D}$.
Given that $\mathcal{B}_{\scriptscriptstyle G}=\mathcal{D}$, and with the application of Lemma 3.4, we derive Equation 5. The proof of Equation 6 follows by identifying the member sets of both families that include $b$ as an element. Similarly, Equation 7 can be established using analogous reasoning. ∎

### v4

@@ 3 Some New Graph Results
Also note that $D$ is a stable set, and since $f^{-1}(D)$ is a subset of $D$, it is also a stable set. Suppose, for contradiction, that $f^{-1}(D)$ is not maximal stable in $H\setminus\Gamma$. This implies there exists a vertex $v\in V(H\setminus\Gamma)$, with $v\notin f^{-1}(D)$, such that the set $\{v\}\cup f^{-1}(D)$ is stable in $H\setminus\Gamma$. Since all vertices in $\Gamma$ are 2-layered, each vertex in $N_{\scriptscriptstyle H}(\Gamma)$ has an adjacent pendant vertex. Due to the maximality of the stable sets, these pendant vertices must be included in all member sets of $\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$. Consequently, they are present in every member set of $f^{-1}(\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma}))$. Thus, $v\notin N_{\scriptscriptstyle H}(\Gamma)$. Therefore, $\{v\}\cup D=\{v\}\cup f^{-1}(D)\cup\Gamma$ would be stable in $H$, which contradicts the assumption that $D$ is a maximal stable set in $H$. Hence, $f^{-1}(D)\in\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})$ and the mapping is well-defined.
It is trivial that $f$ is a bijective map. The one-to-one correspondence between the families $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})$ and $\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ establishes Equation 1. Now note that $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})$ and $\mathcal{B}_{\scriptscriptstyle H}(\{a\}\cup\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ are precisely the subfamilies of $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})$ and $\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ respectively, obtained by restricting to those member sets that contain the vertex $a$. Observe that the bijection defined by $f(B)=B\cup\Gamma$ does not add or remove any $a\in V(H)\setminus\Gamma$; that is, it does not affect membership of $a$. Hence, $f$ restricts to a bijection from $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})$ to $\mathcal{B}_{\scriptscriptstyle H}(\{a\}\cup\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$. This justifies Equation 2. ∎
The Figure 4 gives an illustration of Lemma 3.4, showing the partition of $[n]$ into $\Theta$ and $\Theta^{\mathsf{c}}$, with the subset $\Gamma\subseteq\Theta^{\mathsf{c}}$. Only the 2-layered vertices (in blue) are drawn, while the remaining vertices and edges are omitted. The graph $H$ is shown in green, and the partition $H\setminus\Gamma$ is indicated in purple.
@@
Let $H$ be a graph, and suppose $n]\subseteq V(H)$ contains no isolated vertices. For any $\Theta\subseteq[n]$, $\Gamma_{\scriptscriptstyle 1},\Gamma_{\scriptscriptstyle 2}\subseteq\Theta^{\mathsf{c}}$ and $a\in V(H)$, if $\Gamma_{\scriptscriptstyle 1}\neq\Gamma_{\scriptscriptstyle 2}$, then,
$\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 1}}))\cap\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 2}}))=\emptyset$ (3)
$\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 1}}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 1}}))\cap\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 2}}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 2}}))=\emptyset$ (4)
@@
Consider two distinct subsets $\Gamma_{\scriptscriptstyle 1}$ and $\Gamma_{\scriptscriptstyle 2}$ of $\Theta^{\mathsf{c}}$. Without loss of generality, assume that there exists an element $m\in\Gamma_{\scriptscriptstyle 1}$ such that $m\notin\Gamma_{\scriptscriptstyle 2}$. Suppose $B$ is an element of both $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 1})})$ and $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 2})})$. Since $m\in\Gamma_{\scriptscriptstyle 1}$, the maximal stable sets in $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 1})})$ avoid all vertices in $N_{\scriptscriptstyle H}(m)$. Therefore, $B\cap N_{\scriptscriptstyle H}(m)=\emptyset$. On the other hand, since $B$ is also an element of $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 2})})$, it must be a maximal stable set in $H\setminus\Gamma_{\scriptscriptstyle 2}$ that excludes vertices from $\Theta^{\mathsf{c}}$. Given that $m\in\Theta^{\mathsf{c}}$, the maximality of $B$ implies that $B\cap N_{\scriptscriptstyle H}(m)\neq\emptyset$.
This contradiction shows that $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 1}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 1})})\cap\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 2}}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 2})})$ must be empty. Furthermore, it follows that $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})\subseteq\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})$. Consequently, we deduce that $\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 1}}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 1})})\cap\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma_{\scriptscriptstyle 2}}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma_{\scriptscriptstyle 2})})=\emptyset$. ∎
@@
Let $G$ be a graph with a decomposition $\{H,I\}$ such that $V(H)\cap V(I)=[n]$, where $[n]$ is a stable set and no vertex of $[n]$ is isolated in either $H$ or $I$. For $a\in V(I)\setminus[n]$ and $b\in[n]$, we have,
$w_{\scriptscriptstyle G}=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle I\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (5)
$w_{\scriptscriptstyle G}(a)=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle I\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (6)
$w_{\scriptscriptstyle G}(b)=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle I\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (7)
@@
$\mathcal{B}_{\scriptscriptstyle G}=\left\{B_{1}\cup B_{2}\ \middle|\begin{array}[]{l}B_{1}\in\mathcal{B}_{\scriptscriptstyle I\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)}),B_{2}\in\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\\ \Theta\subseteq[n],\hskip 14.22636pt\Gamma\subseteq\Theta^{\mathsf{c}},\hskip 14.22636pt\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma\end{array}\right\}.$
@@
For simplicity, let us denote the family on the right-hand side by $\mathcal{D}$. Consider any member set $B_{1}\cup B_{2}$ in $\mathcal{D}$. It is straightforward to observe that $B_{1}\cup B_{2}$ forms a stable set in $G$. If this set is not maximal in $G$, then there must exist a vertex $v\in V(G)$ such that $\{v\}\cup B_{1}\cup B_{2}$ is also a stable set. However, due to the maximality of $B_{1}$ and $B_{2}$ in the subgraphs $I\setminus\Gamma$ and $H\setminus\Psi$, respectively, the vertex $v$ cannot belong to either of these subgraphs. Since $I\setminus\Gamma\cup H\setminus\Psi=G$, there is no such vertex $v$ in the graph $G$. Therefore, $B_{1}\cup B_{2}$ must be a maximal stable set in $G$, implying that $B_{1}\cup B_{2}\in\mathcal{B}_{\scriptscriptstyle G}$ and hence $\mathcal{D}\subseteq\mathcal{B}_{\scriptscriptstyle G}$.
@@
Now, consider any $B\in\mathcal{B}_{\scriptscriptstyle G}$. Define the subsets $\Theta\subseteq[n]$ and $\Theta^{\mathsf{c}}=[n]\setminus\Theta$ by setting $\Theta=B\cap[n]$. Next, define $\Gamma,\Psi\subseteq\Theta^{\mathsf{c}}$ as the sets of vertices in $[n]$ such that $N_{\scriptscriptstyle I}(\Gamma)\cap B=\emptyset$ and $N_{\scriptscriptstyle H}(\Psi)\cap B=\emptyset$, respectively. It is evident that $\Gamma\cap\Psi=\emptyset$, because if there were a vertex $v\in\Theta^{\mathsf{c}}$ with $B\cap N_{\scriptscriptstyle G}(v)=B\cap(N_{\scriptscriptstyle I}(v)\cup N_{\scriptscriptstyle H}(v))=(B\cap N_{\scriptscriptstyle I}(v))\cup(B\cap N_{\scriptscriptstyle H}(v))=\emptyset$, this would contradict the maximality of $B$ as a stable set [^4] Since $v$ and none of its neighbors would be included in $B$. Furthermore, $B\cap(I\setminus\Gamma)$ contains all vertices in $\Theta$ but excludes vertices from both $\Theta^{\mathsf{c}}$ and $N_{\scriptscriptstyle I}(\Gamma)$. Notably, this set is stable since it is a subset of $B$.
@@
Assume that $B$ is not maximal in $I\setminus\Gamma$, and there exists a vertex $v\in V(I\setminus\Gamma)$ such that $v\notin B$ and $(\{v\}\cup B)\cap(I\setminus\Gamma)$ is a stable set. For each $u\in\Theta^{\mathsf{c}}\setminus\Gamma$, by definition, $N_{\scriptscriptstyle I}(u)\cap B\neq\emptyset$, which implies that $v$ cannot belong to $\Theta^{\mathsf{c}}\setminus\Gamma$. Consequently, $v\notin[n]$. Note that $v$ also does not belong to $N_{\scriptscriptstyle I}(\Theta)$, as $(\{v\}\cup B)\cap(I\setminus\Gamma)$ remains stable. From this, one can observe that $(\{v\}\cup B)\cap(H\setminus\Psi)$ is also stable, and consequently, $B\cup\{v\}=((\{v\}\cup B)\cap(I\setminus\Gamma))\cup((\{v\}\cup B)\cap(H\setminus\Psi))$ forms a stable set in $G$, which contradicts the maximality of $B$. Therefore, $B\cap(I\setminus\Gamma)\in\mathcal{B}_{\scriptscriptstyle I\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})$. Similarly, one can show that $B\cap(H\setminus\Psi)\in\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})$, and hence $B\in\mathcal{D}$. This implies that $\mathcal{B}_{\scriptscriptstyle G}\subseteq\mathcal{D}$.
With the application of Lemma 3.5, one can see that any $B\in\mathcal{D}$ has a unique $B_{1}$ and $B_{2}$ associated where $B_{1}\cup B_{2}=B$, with $B_{1}\in\mathcal{B}_{\scriptscriptstyle I\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})$ and $B_{2}\in\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})$. This, together with $\mathcal{B}_{\scriptscriptstyle G}=\mathcal{D}$, derives Equation 5. Now notice that for any $a\in V(I)\setminus[n]$, the family $\mathcal{B}_{\scriptscriptstyle G}(a)$ consists of the maximal stable sets of $G$ containing $a$. Due to the one-to-one correspondence with $\mathcal{D}$, this must be the collection of stable sets in $\mathcal{D}$ that contain $a$. That is,
$\mathcal{B}_{\scriptscriptstyle G}(a)=\left\{B_{1}\cup B_{2}\ \middle|\begin{array}[]{l}B_{1}\in\mathcal{B}_{\scriptscriptstyle I\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)}),B_{2}\in\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\\ \Theta\subseteq[n],\hskip 14.22636pt\Gamma\subseteq\Theta^{\mathsf{c}},\hskip 14.22636pt\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma\end{array}\right\}.$
Thus, Equation 6 is established. Now, for Equation 7, consider any $b\in[n]$. One can see that $\mathcal{B}_{\scriptscriptstyle G}(b)$ is nothing but the collection of member sets of $\mathcal{D}$ which contain $b$. As $b\in V(H)\cap V(I)$, it directly follows that
$\mathcal{B}_{\scriptscriptstyle G}(b)=\left\{B_{1}\cup B_{2}\ \middle|\begin{array}[]{l}B_{1}\in\mathcal{B}_{\scriptscriptstyle I\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)}),B_{2}\in\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\\ \Theta\subseteq[n],\hskip 14.22636pt\Gamma\subseteq\Theta^{\mathsf{c}},\hskip 14.22636pt\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma\end{array}\right\}.$
An illustration of Lemma 3.6 is provided in Figure 4. In this figure, the graph $G$ is decomposed into two subgraphs $H$ and $I$, and the set of common vertices $[n]=[4]$ is partitioned into $\Theta=\{1\}$ and $\Theta^{\mathsf{c}}=\{2,3,4\}$. Furthermore, $\Theta^{\mathsf{c}}$ is divided into the disjoint subsets $\Gamma=\{2\}$ and $\Psi=\{4\}$, as defined in the lemma. Consider the maximal stable set $\{1,6,9,11,13,14,17,18\}\in\mathcal{B}_{\scriptscriptstyle G}(\Theta,\overline{\Theta^{\mathsf{c}}})$. Observe that this set can be expressed as the union of $\{1,6,9,17,18\}$ and $\{1,11,13,14\}$. These two sets belong to the families $\mathcal{B}_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})$ and $\mathcal{B}_{\scriptscriptstyle I\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})$, respectively. This demonstrates the bijection described in the proof of Lemma 3.6.